# Crushing, Comminution and Fracture: Extreme Particle Deformation in Three-Dimensional Granular Aggregates

### v1

@@ Abstract
 [^1] Authors are listed in alphabetical order. [^2] Authors are listed in alphabetical order. [^3] Authors are listed in alphabetical order.
@@ 5 Self-contact Forces
Repulsive Contact Bond Force Short-range repulsive bond force between particles, preventing overlap. $\mathbf{f}_{r}(\mathbf{y},\mathbf{x},t)=\begin{cases}-K_{n}(R_{c}-|\mathbf{p}(\mathbf{y},t)-\mathbf{p}(\mathbf{x},t)|)V_{x}V_{y}e(\mathbf{y},\mathbf{x},t),&\text{if }|\mathbf{p}(\mathbf{y},t)-\mathbf{p}(\mathbf{x},t)|<R_{c},\\ 0,&\text{otherwise}.\end{cases}$
Nonlocal Friction Bond Force Tangential friction bond force, modeling sliding resistance between particles. $\mathbf{f}_{f}(\mathbf{y},\mathbf{x},t)=\begin{cases}-\mu|\mathbf{f}_{r}(\mathbf{y},\mathbf{x},t)|\mathbf{e}_{\perp}(\mathbf{y},\mathbf{x},t),&\text{if}|\mathbf{v}_{\perp}(\mathbf{y},\mathbf{x},t)|>v_{\text{thr}},\\ -\frac{1}{\epsilon_{f}}|\mathbf{v}_{\perp}(\mathbf{y},\mathbf{x},t)|\mathbf{e}_{\perp}(\mathbf{y},\mathbf{x},t),&\text{otherwise}.\end{cases}$
Normal Damping Bond Force Bond Energy dissipation during particle interactions. $\mathbf{f}_{d}(\mathbf{y},\mathbf{x},t)=\begin{cases}-\beta_{d}\left(\frac{\mathbf{v}(\mathbf{y},\mathbf{x},t)}{|\mathbf{v}(\mathbf{y},\mathbf{x},t)|}\cdot e(\mathbf{y},\mathbf{x},t)\right)e(\mathbf{y},\mathbf{x},t),&\text{if }|\mathbf{p}(\mathbf{y},t)-\mathbf{p}(\mathbf{x},t)|<R_{c},\\ 0,&\text{otherwise}.\end{cases}$
@@
Body Force Density (e.g., Gravity) External body force densities acting on all particles, such as gravity. $\mathbf{b}(x,t)$
Wall Repulsive Force Density Force density that prevents particles from penetrating container walls. $\mathbf{F}_{r}^{\text{wall}}(\mathbf{x},t)=\begin{cases}-K_{n}(R_{c}-|\mathbf{c}_{l}(\mathbf{x},t)-\mathbf{p}(\mathbf{x},t)|)|S_{l}(\mathbf{x},t)|e_{l}(\mathbf{x},t),&\text{if }d_{l}(\mathbf{x},t)<R_{c},\\ 0,&\text{otherwise}.\end{cases}$
Wall Friction Force Density Tangential friction force density between particles and container walls. $\mathbf{F}_{f}^{\text{wall}}(\mathbf{x},t)=-\mu|\mathbf{F}_{r}^{\text{wall}}(\mathbf{x},t)|\mathbf{e}_{\perp}(\mathbf{x},t)$
@@
Self-Repulsive Bond Force Repulsive bond force between parts of the same particle to avoid overlap. $\mathbf{f}_{r}^{\text{self}}(\mathbf{x},\mathbf{y},t)=\begin{cases}c_{w}\left(\frac{|\mathbf{\xi}+\eta|-|\mathbf{\xi}|}{|\mathbf{\xi}|}\right),&\text{if }|\mathbf{\xi}|<R_{c},\\ c_{w}\frac{|\mathbf{\xi}+\eta|-R_{c}}{R_{c}},&\text{if }R_{c}<|\mathbf{\xi}|<\epsilon.\end{cases}$
Table 2: Categorization of Forces in the PeriDEM Particle Dynamics Model. All bond forces are in units of force per unit volume squared.
@@ 10.1 Acknowledgements
This material is based upon work supported by the U. S. Army Research Laboratory and the U. S. Army Research Office under Contract/Grant Number W911NF-19-1-0245. Portions of this research were conducted with high performance computing resources provided by Louisiana State University (http://www.hpc.lsu.edu).

### v2

@@ Abstract
journal: Computer Methods in Applied Mechanics and Engineering [^1] Authors are listed in alphabetical order. [^2] Authors are listed in alphabetical order. [^3] Authors are listed in alphabetical order.
@@ 5 Self-contact Forces
Repulsive Contact Force Short-range repulsive bond force preventing particle overlap $\mathbf{f}_{r}(\mathbf{y},\mathbf{x},t)=\begin{cases}-K_{n}(R_{c}-|\boldsymbol{\delta}_{yx}|)\,\mathbf{e},&\text{if }|\boldsymbol{\delta}_{yx}|<R_{c}\\ 0,&\text{otherwise}\end{cases}$
Nonlocal Stick Slip Friction Tangential friction force modeling sliding resistance $\mathbf{f}_{f}(\mathbf{y},\mathbf{x},t)=\begin{cases}-\mu|\mathbf{f}_{r}|\,\mathbf{e}_{\perp},&\text{if }|\mathbf{v}_{\perp}|>v_{\text{thr}}\\ \displaystyle-\frac{1}{\epsilon_{f}}|\mathbf{v}_{\perp}|\,\mathbf{e}_{\perp},&\text{otherwise}\end{cases}$
Normal Damping Force Energy dissipation during interaction $\mathbf{f}_{d}(\mathbf{y},\mathbf{x},t)=\begin{cases}-\beta_{d}\left(\dfrac{\mathbf{v}}{|\mathbf{v}|}\cdot\mathbf{e}\right)\mathbf{e},&\text{if }|\boldsymbol{\delta}_{yx}|<R_{c}\\ 0,&\text{otherwise}\end{cases}$
@@
Body Force Density (e.g., Gravity) External force acting on all particles $\mathbf{b}(\mathbf{x},t)$
Wall Repulsive Force Density Prevents penetration into container walls $\mathbf{F}_{r}^{\text{wall}}(\mathbf{x},t)=\begin{cases}-K_{n}(R_{c}-|\mathbf{c}_{l}-\mathbf{p}|)\,|S_{l}|\,\mathbf{e}_{l},&\text{if }d_{l}(\mathbf{x},t)<R_{c}\\ 0,&\text{otherwise}\end{cases}$
@@
Self-Repulsive Bond Force Avoids overlap within the same particle $\scriptsize\mathbf{f}_{r}^{\text{self}}(\mathbf{x},\mathbf{y},t)=\begin{cases}\displaystyle c_{w}\frac{|\boldsymbol{\delta}_{xy}|-|\mathbf{y}-\mathbf{x}|}{|\mathbf{y}-\mathbf{x}|}\mathbf{e}(\mathbf{y},\mathbf{x},t)\chi_{\{|\boldsymbol{\delta}_{xy}|<|\mathbf{y}-\mathbf{x}|\}},&|\mathbf{y}-\mathbf{x}|<R_{c}\\ \displaystyle c_{w}\frac{|\boldsymbol{\delta}_{xy}|-R_{c}}{R_{c}}\mathbf{e}(\mathbf{y},\mathbf{x},t)\chi_{\{|\boldsymbol{\delta}_{xy}|<R_{c}\}},&\text{otherwise}\end{cases}$
Table 2: Categorization of force terms in the PeriDEM particle dynamics model. All bond forces are expressed per unit volume squared. Here, $\boldsymbol{\delta}_{yx}=\mathbf{p}(\mathbf{y},t)-\mathbf{p}(\mathbf{x},t)$ and $\boldsymbol{\delta}_{xy}=\mathbf{p}(\mathbf{x},t)-\mathbf{p}(\mathbf{y},t)$ denote the relative positions used in bond evaluations.
@@ 10.1 Acknowledgements
This material is based upon work supported by the U. S. Army Research Laboratory and the U. S. Army Research Office under Contract/Grant Number W911NF-19-1-0245 and W911NF-24- 2-0184. Portions of this research were conducted with high performance computing resources provided by Louisiana State University (http://www.hpc.lsu.edu).
Data availability The code used to generate the results in this study is available at
https://github.com/DavoodDamircheli/PeriDEM.